Verify that 𝑦1(𝑥) = 1 and 𝑦2(𝑥) = 𝑥^(1/2) are solutions of the differential equation 𝑦𝑦′′ + (𝑦′)^2 = 0 for 𝑥 > 0. Then show that 𝑦 = 𝑐1 + 𝑐2𝑥^(1/2) is not, in general, a solution to the equation. Explain why this does not contradict superposition principle.


1
Expert's answer
2022-04-27T01:22:27-0400

Compute derivatives of "y_1(x)" : "(y_1(x))'=0", "(y_1(x))''=0." Thus, we get: "y_1(x)(y_1(x))''+(y_1(x)')^2=0". Compute derivatives of "y_2(x)": "(y_2(x))'=\frac12x^{-\frac12}", "(y_2(x))''=-\frac14x^{-\frac32}". We get: "y_2(x)(y_2(x))''+(y_2(x)')^2=0".

Consider the function "y=c_1+c_2x^{\frac12}". Compute the derivatives: "y'=\frac12c_2x^{-\frac12}", "y''=-\frac14c_2x^{-\frac32}". Consider the expression: "y(x)(y(x))''+(y')^2=-\frac14c_1c_2x^{-\frac32}-\frac14c_2^2x^{-1}+\frac14c_2^2x^{-1}=-\frac14c_1c_2x^{-\frac32}"

As we can see, "y(x)" is a solution if and only if either "c_1=0" or "c_2=0". It means that in general "y(x)" is not the solution. It is due to the fact that the equation "yy''+(y')^2=0" is nonlinear. Linear superposition principle holds only for linear differential equations.